OpenStudy (anonymous):

PLEASE HELP!!!!! MEDAL WILL BE REWARDED!!!! Find an equation in standard form for the ellipse with the vertical major axis of length 18, and minor axis of length 6.

OpenStudy (anonymous):

where is it centered

OpenStudy (anonymous):

b=3 a=4 if centered at origin use \[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\]